An open water tank has an orifice 10cm below the surface of water .if there is no loss of energy ,then what is the speed of efflux ???????????????????????????????? (take g=10 m/s square )
1
Expert's answer
2013-11-28T14:05:18-0500
v^2/2 + gz + p/rho = const.

we have the same pressure, as the tank is opened, hence

v^2/2+gh=const.

Hence

v^2/2 = gz

v = sqrt(2gz) = sqrt(2*10*0.1) = sqrt(2) = 1.41 m/s